OpenStudy (anonymous):

solve this problem 8(5a+3)=64

OpenStudy (anonymous):

go to wolframalpha.com its a calculator website for math and it will show you the steps too. In this one a=1

OpenStudy (anonymous):

8(5a+3)=64 40a+24=64 -24 -24 40a=40 40a divided by 40 a =1